# The Cosmological CPT Theorem

### v2

@@ 6.2 Perturbative Discrete Symmetries
where, apart from the overall factor of $\sqrt{g}$, only integer powers of the metric appear in $\mathcal{L}$ [^43] For spinor fields, the propagator term contains a factor of $\gamma_{\mu}$ which scales like the square-root of a metric. So in theories with fermions, we expect an additional power of (1)(-1)( - 1 ) in $\bf{CRT}$ for each fermion propagator, in any dimension $D$. In this section, we are defining $I$ with the conventions of the Euclidean action, so that the amplitude always goes like $e^{-I}$, even in Lorentzian signature; thus the Lorentzian action is imaginary.
@@
On the other hand, when $D=d+1$ is odd, the Lagrangian changes sign: $\mathcal{L}\to-\mathcal{L}$. In any Feynman-Witten diagram, this provides an extra minus sign for every vertex $V$, and also every internal edge $I$ (because the sign of the propagator term also reverses). In any connected diagram, the Euler formula $V-I=1-L$ allows us to write the change of any unrenormalised Feynman-Witten amplitude ${\cal A}$ in terms of the number of loops $L$:
@@
where $\mathcal{A}^{(L)}$ denotes a perturbative amplitude computed at some loop order $L$ (e.g. $L=0$ corresponds to tree-level and $L=1$ corresponds to 111-loop order), and we allow $d$ to be non-integer. The reason we keep saying “unrenormalised”, is that this pattern does not continue to hold when doing renormalisation theory. The reason is that the UV cutoff $\epsilon$ has units of length, and so it implicitly depends on the value of $g_{ab}$ and hence $\Omega$. If, for example, one has a $\log$ divergence of the form $C\log(\epsilon)$, then when $\Omega\to e^{-i\pi}\Omega$, this will also send $\epsilon\to e^{-i\pi}\epsilon$, thus producing the shift:
@@
$\mathbf{CRT}:\qquad\psi_{n}^{(L)*}(\eta;y_{1},\dots,y_{d})\;=$ $\!\!e^{i\pi(d+1)(L-1)}$ $\!\!\!\!\!\psi_{n}^{(L)}(\,e^{-i\pi}\eta;-y_{1},\dots,-y_{d})\,;$ (6.25)
$\mathbf{D}_{-1}^{\pm}:\qquad\psi_{n}^{(L)}(\eta;y_{1},\dots,y_{d})\;=$ $\!\!\!\!\!\!\psi_{n}^{(L)}(e^{\pm i\pi}\eta;e^{\pm i\pi}y_{1},\dots,e^{\pm i\pi}y_{d})\,;$ (6.26)
$\mathbf{RR}:\qquad\psi_{n}^{(L)*}(\eta;y_{1},\dots,y_{d})\;=$ $\!\!e^{i\pi(d+1)(L-1)}$ $\!\!\!\!\!\psi_{n}^{(L)}(\;\;\eta\;;-e^{i\pi}y_{1},\dots,-e^{i\pi}y_{d})\,,$ (6.27)
@@
$\mathbf{CRT}:\qquad\psi_{n}^{(L)*}(\eta;k_{1},\dots,k_{d})\;=$ $\!\!e^{i\pi(d+1)(L-1)}$ $\!\!\!\!\!\!\psi_{n}^{(L)}(\,e^{-i\pi}\eta;k_{1},\dots,k_{d})\,;$ (6.28)
$\mathbf{D}_{-1}^{\pm}:\qquad\psi_{n}^{(L)}(\eta;k_{1},\dots,k_{d})\;=$ $\!\!\qquad\,\,e^{\pm i\pi d}$ $\!\!\!\!\!\!\psi_{n}^{(L)}(e^{\pm i\pi}\eta;e^{\mp i\pi}k_{1},\dots,e^{\mp i\pi}k_{d})\,;$ (6.29)
$\mathbf{RR}:\qquad\psi_{n}^{(L)*}(\eta;k_{1},\dots,k_{d})\;=$ $\;\;e^{i\pi((d+1)L-1)}$ $\!\!\!\!\!\psi_{n}^{(L)}(\,\eta\,;e^{-i\pi}k_{1},\dots,e^{-i\pi}k_{d})\,.$ (6.30)
As this section considers only the case of a parity-even Lagrangian, technically the arguments in this section don’t care about whether we reflect the $y$-coordinates in $\bf{CRT}$ or $\bf{RR}$. However, we have put the signs in the correct place for parity-odd theories as well, as can be deduced in Section 6.3 where we do not restrict to parity-even theories. Also, we are assuming throughout this paper that none of the fields involved are in the principal series. [^47] This restriction applies here because of a subtle caveat to our arguments above, that the Weyl factor only affects the $\sqrt{g}$ factor of the metric. Specifically, we have assumed that we are doing perturbation theory in $m^{2}$ as well as the non-linear interactions. However, in QFT it is more common to write the propagator as $1/(p^{2}-m^{2})$ and calculate the answer non-perturbatively in $m^{2}$. When you do it this way, one sometimes finds effects that are non-perturbative in $m^{2}$. (E.g. in flat $D=3$ dimensional space, the massive propagator is given by the Yukawa form $e^{-mr}/r$, which is non-analytic in $m^{2}$.) This phenomenon arises because there is a branch cut in the amplitude as a function of $m^{2}$, which is related to a change of boundary conditions out at infinity. As a result the $\Omega\to e^{-i\pi}\Omega$ rotation cannot be summarised only by the $\sqrt{g}$ rotation. We believe that in de Sitter, this non-analyticity exists only when at least one field line (external or internal) is in the principal series (so that $\Delta\neq\Delta^{*}$).
@@ 6.3 Cosmological Optical Theorem for Flat FLRW
where we define each permutation symbol so that it is invariant under all real coordinate transforms in (time / space) except those for which $\det$ Jacobian =1=-1= - 1, in which case the (time / space) permutation symbol flips sign. [^48] If a Lorentz covariant parity-odd theory is desired, one simply requires the epsilon symbols to come together in the form of the usual spacetime permutation symbol: $\bm{\epsilon}^{a_{0}a_{1}\ldots a_{d}}=(d+1)\bm{\epsilon}^{[a_{0}}\bm{\epsilon}^{a_{1}a_{2}\ldots a_{d}]}$. However, for any analytic continuation that is continuously connected to the identity, the $\bm{\epsilon}$’s have to remain the same sign throughout, as the analytic continuation of 1 is 1. Now let a given amplitude be labelled $T=\pm 1$ according to whether it depends on $\bm{\epsilon}^{t}$ an (even/odd) number of times, and let it be labelled $X=\pm 1$ according to whether it depends on $\bm{\epsilon}^{i_{1}i_{2}\ldots i_{d}}$ an (even/odd) number of times.
@@
$\psi_{n}^{*}(\eta;y_{1},\dots,y_{d};\Omega)=T\psi_{n}(\eta;e^{i\pi}y_{1},\dots,e^{i\pi}y_{d};e^{-i\pi}\Omega)\,,$ (6.33)
@@
$\psi_{n}(\eta;-y_{1},\dots,-y_{d};\Omega)=X^{d}\psi_{n}(\eta;y_{1},\dots,y_{d};\Omega)\,.$ (6.34)
@@
$\psi_{n}^{(L)}(\eta;y_{1},\dots,y_{d};e^{i\pi}\Omega)=TX^{d}(-1)^{D(L-1)}\psi_{n}^{(L)}(\eta;y_{1},\dots,y_{d};\Omega)=TX^{d}e^{i\pi D(L-1)}\psi_{n}^{(L)}(\eta;y_{1},\dots,y_{d};\Omega)\,.$ (6.35)
@@
$\mathbf{RR}:\qquad\psi_{n}^{(L)*}(\eta;y_{1},\dots,y_{d})=e^{i\pi(d+1)(L-1)}\psi_{n}^{(L)}(\eta;-e^{i\pi}y_{1},\dots,-e^{i\pi}y_{d})\,.$ (6.36)
@@ 7.1 Unitarity and the Bunch-Davies Vacuum
$\rightarrow e^{-i\theta}\textbf{k}\,.$ k . (7.11)
@@
This is a linear combination of sines and cosines and so could equivalently be represented as a shifted sine function which is non-zero at $\theta=0$. Therefore, this function is guaranteed to change sign over the range $0\leq\theta\leq\pi$ (or $-\pi\leq\theta\leq 0$) and we must also rotate $\varepsilon$ alongside $\eta$, $\varepsilon\rightarrow\varepsilon e^{i\theta}$, so that the exponent is
@@ 7.2 The Loop Momentum Integral
where the prime in $\psi^{\prime(L)}_{n}$ is used to explicitly highlight that these wavefunction coefficients contain the $\delta$-function whilst those without the $\delta$-function will be denoted by $\psi_{n}^{(L)}$, and we have used the shorthand notation $\psi^{(L)}_{n}(\eta_{0};\textbf{k};\varepsilon)\equiv\psi_{n}^{(L)}(\eta;k_{1},\dots,k_{d};\epsilon)$. The external energies are indicated by $k_{a}=\lvert\textbf{k}_{a}\rvert$ whilst the $p_{i}$ are the internal energies which are built from the $\textbf{k}_{a}$ plus potentially some of the loop momenta, $\textbf{P}_{l}$. Note that here we have made the $\varepsilon$ prescription explicit and that there are no derivatives acting on any of our propagators. This is purely for the sake of notational simplicity and adding any derivatives will not add any technical complications.
@@
The first component we will look at is the time integral which runs from $\infty$- ∞ to $\eta_{0}$ when we have a transformation involving the time coordinate it is these limits that pick up a minus sign. To remove this we redefine the variables inside the integral with $\eta_{v}\rightarrow-\eta_{v}$ [^52] Note that whether this minus sign is $e^{i\pi}$ or $e^{-i\pi}$ will be important but, as it can rotate in a different way depending on which transformation we are considering, we will keep things vague at this point. When we write explicit expressions for the transformations we will be careful to make clear in which direction the rotation goes.. The effect of this on the measure factor is then straightforward as everything is just some power of $\eta$.
@@
For positive $\lambda$ we can see that the scaling on each side agrees. Naively, on the left hand side continuing $\lambda\rightarrow-\lambda$ would appear to have no impact. However, from the right hand side we can see how our analytic continuation in the energies alters this conclusion, instead recovering a factor of $(-1)^{d}$. [^53] Now that we have identified how to analytically continue loop energies, this removes the barrier which prevented the authors of [28] from deriving the cosmological equivalent of cutting rules from generalised Unitarity (see [130] for a pedagogical review of flat space generalised Unitarity). To see this, we first note that the region, $\Gamma$, depends only on the squares of the energies and so is unchanged. However, each of our square roots of energies does pick up a factor of 1-1- 1. As we now have a potentially non-integer exponent in the dimension of the integral $d$ we will need to be careful with $e^{i\pi d}$ vs $e^{-i\pi d}$. For $\bf{RR}$, in accordance with the discussion in Section 7.1, all rotations in the energy must be taken in the lower half plane and so $\lambda=\lambda e^{-i\pi}$. Therefore we have,
@@
$\lim_{\eta\rightarrow\eta_{0}}K_{k}(\eta)$ =1=1\,,= 1 , $\lim_{\eta\rightarrow-\infty}K_{k}(\eta,\eta^{\prime})$ $\propto e^{ik\eta}\,,$ (7.24)
$\lim_{\eta,\eta^{\prime}\rightarrow\eta_{0}}G_{k}(\eta,\eta^{\prime})$ =0=0\,,= 0 , $\lim_{\eta,\eta^{\prime}\rightarrow-\infty}G_{k}(\eta,\eta^{\prime})$ $\propto e^{ik\eta}\,.$ (7.25)
@@
$\mathcal{O}_{\textbf{k}}^{*}(\eta)K_{k}^{*}(\eta)$ =0=0,= 0 , $\lim_{\eta\rightarrow\eta_{0}}K_{k}^{*}(\eta)$ =1=1\,,= 1 , $\lim_{\eta\rightarrow-\infty}K_{k}^{*}(\eta)$ $\propto e^{-ik\eta}\,,$ (7.29)
$\mathcal{O}_{e^{-i\pi}\textbf{k}}(\eta)K_{e^{-i\pi}k}(\eta)$ =0=0,= 0 , $\lim_{\eta\rightarrow\eta_{0}}K_{e^{-i\pi}k}(\eta)$ =1=1\,,= 1 , $\lim_{\eta\rightarrow-\infty}K_{e^{-i\pi}k}(\eta)$ $\propto e^{-ik\eta}\,.$ (7.30)
@@
$-\mathcal{O}_{\textbf{k}}^{*}(\eta)G_{k}^{*}(\eta,\eta^{\prime})$ $=-i\delta(\eta-\eta^{\prime}),$ $\lim_{\eta\rightarrow\eta_{0}}-G_{k}^{*}(\eta,\eta^{\prime})$ =0=0\,,= 0 , $\lim_{\eta\rightarrow-\infty}-G_{k}^{*}(\eta,\eta^{\prime})$ $\propto e^{-ik\eta}\,,$ (7.32)
$\mathcal{O}_{e^{-i\pi}\textbf{k}}(\eta)G_{e^{-i\pi}k}(\eta,\eta^{\prime})$ $=-i\delta(\eta-\eta^{\prime}),$ $\lim_{\eta\rightarrow\eta_{0}}G_{e^{-i\pi}k}(\eta,\eta^{\prime})$ =0=0\,,= 0 , $\lim_{\eta\rightarrow-\infty}G_{e^{-i\pi}k}(\eta,\eta^{\prime})$ $\propto e^{-ik\eta}\,,$ (7.33)
@@
$\mathcal{O}_{\textbf{k}}^{*}(\eta)K_{k}^{*}(\eta)$ =0=0,= 0 , $\lim_{\eta\rightarrow\eta_{0}}K_{k}^{*}(\eta)$ =1=1\,,= 1 , $\lim_{\eta\rightarrow-\infty}K_{k}^{*}(\eta)$ $\propto e^{-ik\eta}\,,$ (7.35)
$\mathcal{O}_{\textbf{k}}(e^{-i\pi}\eta)K_{k}(e^{-i\pi}\eta)$ =0=0,= 0 , $\lim_{\eta\rightarrow\eta_{0}}K_{k}(e^{-i\pi}\eta)$ =1=1\,,= 1 , $\lim_{\eta\rightarrow-\infty}K_{k}(e^{-i\pi}\eta)$ $\propto e^{-ik\eta}\,.$ (7.36)
@@
$\lim_{\eta\rightarrow\eta_{0}}G_{k}^{*}(\eta,\eta^{\prime})$ =0=0\,,= 0 , $\lim_{\eta\rightarrow\eta_{0}}e^{i\pi(d+1)}G_{k}(e^{-i\pi}\eta,e^{-i\pi}\eta^{\prime})$ =0=0\,,= 0 , (7.39)
@@
[…]
@@
where in the penultimate line we have employed the Feynman-Euler relationship, $V-I+L=1$ and the rotation on k applies to all internal and external energies. Next we investigate Scale Invariance which rotates all momenta magnitudes as well as the conformal time,
@@
×$\qquad\qquad\qquad\qquad\quad\times\prod_{a}^{n}K_{e^{\mp i\pi}k_{a}}(e^{\pm i\pi}\eta)\prod_{i}^{I}G_{e^{\mp i\pi}p_{i}}(e^{\pm i\pi}\eta,e^{\pm i\pi}\eta^{\prime})\delta^{d}\left(e^{\mp i\pi}\sum_{a}^{n}\textbf{k}_{a}\right)\,$ (7.52)
@@
[…]

### v3

@@ 6.2 Perturbative Discrete Symmetries
where, apart from the overall factor of $\sqrt{g}$, only integer powers of the metric appear in $\mathcal{L}$. [^43] For spinor fields, the propagator term contains a factor of $\gamma_{\mu}$ which scales like the square-root of a metric. So in theories with fermions, we expect an additional power of (1)(-1) in $\bf{CRT}$ for each fermion propagator, in any dimension $D$. In this section, we are defining $I$ with the conventions of the Euclidean action, so that the amplitude always goes like $e^{-I}$, even in Lorentzian signature; thus the Lorentzian action is imaginary.
@@
On the other hand, when $D=d+1$ is odd, the Lagrangian changes sign: $\mathcal{L}\to$-$\mathcal{L}$. In any Feynman-Witten diagram, this provides an extra minus sign for every vertex $V$, and also every internal edge $I$ (because the sign of the propagator term also reverses). In any connected diagram, the Euler formula VI=1LV-I=1-L allows us to write the change of any unrenormalised Feynman-Witten amplitude $\cal A$} in terms of the number of loops $L$:
@@
where $\mathcal{A}^{(L)}$ denotes a perturbative amplitude computed at some loop order $L$ (e.g. $L=0$ corresponds to tree-level and $L=1$ corresponds to $1$-loop order), and we allow $d$ to be non-integer. The reason we keep saying “unrenormalised”, is that this pattern does not continue to hold when doing renormalisation theory. The reason is that the UV cutoff $\epsilon$ has units of length, and so it implicitly depends on the value of $g_{ab}$ and hence $\Omega$. If, for example, one has a $\log$ divergence of the form $C\log(\epsilon)$, then when $\Omega\to e^{-i\pi}\Omega$, this will also send $\epsilon\to e^{-i\pi}\epsilon$, thus producing the shift:
@@
$\mathbf{CRT}:\qquad\psi_{n}^{(L)*}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n})\;=$ $\!\!e^{i\pi(d+1)(L-1)}$ $\!\!\!\!\!\psi_{n}^{(L)}(\,e^{-i\pi}\eta;-\textbf{y}_{1},\dots,-\textbf{y}_{n})\,;$ (6.25)
$\mathbf{D}_{-1}^{\pm}:\qquad\psi_{n}^{(L)}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n})\;=$ $\!\!\!\!\!\!\psi_{n}^{(L)}(e^{\pm i\pi}\eta;e^{\pm i\pi}\textbf{y}_{1},\dots,e^{\pm i\pi}\textbf{y}_{n})\,;$ (6.26)
$\mathbf{RR}:\qquad\psi_{n}^{(L)*}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n})\;=$ $\!\!e^{i\pi(d+1)(L-1)}$ $\!\!\!\!\!\psi_{n}^{(L)}(\;\;\eta\;;-e^{i\pi}\textbf{y}_{1},\dots,-e^{i\pi}\textbf{y}_{n})\,$, (6.27)
@@
$\mathbf{CRT}:\qquad\psi_{n}^{(L)*}(\eta;\textbf{k}_{1},\dots,\textbf{k}_{n})\;=$ $\!\!e^{i\pi(d+1)(L-1)}$ $\!\!\!\!\!\!\psi_{n}^{(L)}(\,e^{-i\pi}\eta;\textbf{k}_{1},\dots,\textbf{k}_{n})\,;$ (6.28)
$\mathbf{D}_{-1}^{\pm}:\qquad\psi_{n}^{(L)}(\eta;\textbf{k}_{1},\dots,\textbf{k}_{n})\;=$ $\!\!\qquad\,\,e^{\pm i\pi d}$ $\!\!\!\!\!\!\psi_{n}^{(L)}(e^{\pm i\pi}\eta;e^{\mp i\pi}\textbf{k}_{1},\dots,e^{\mp i\pi}\textbf{k}_{n})\,;$ (6.29)
$\mathbf{RR}:\qquad\psi_{n}^{(L)*}(\eta;\textbf{k}_{1},\dots,\textbf{k}_{n})\;=$ $\;\;e^{i\pi((d+1)L-1)}$ $\!\!\!\!\!\psi_{n}^{(L)}(\,\eta\,;e^{-i\pi}\textbf{k}_{1},\dots,e^{-i\pi}\textbf{k}_{n})\,.$ (6.30)
As this section considers only the case of a parity-even Lagrangian, technically the arguments in this section don’t care about whether we reflect the $y$-coordinates in $\bf{CRT}$ or $\bf{RR}$. However, we have put the signs in the correct place for parity-odd theories as well, as can be deduced in Section 6.3 where we do not restrict to parity-even theories. Also, we are assuming throughout this paper that none of the fields involved are in the principal series. [^47] For the principal series, one cannot choose a boundary condition which is both self-adjoint and conformally-invariant. If you sacrifice conformal-invariance, then $\Delta$ is not defined. On the other hand, if you sacrifice self-adjointness, there is a subtle caveat to our arguments above, that the Weyl factor only affects the $\sqrt{g}$ factor of the metric. Specifically, we have assumed that we are doing perturbation theory in $m^{2}$ as well as the non-linear interactions. However, in QFT it is more common to write the propagator as $1/(p^{2}-m^{2})$ and calculate the answer non-perturbatively in $m^{2}$. When you do it this way, one sometimes finds effects that are non-perturbative in $m^{2}$. (E.g. in flat $D=3$ dimensional space, the massive propagator is given by the Yukawa form $e^{-mr}/r$, which is non-analytic in $m^{2}$.) This phenomenon arises because there is a branch cut in the amplitude as a function of $m^{2}$, which is related to a change of boundary conditions out at infinity. As a result the $\Omega\to e^{-i\pi}\Omega$ rotation cannot be summarised only by the $\sqrt{g}$ rotation. We believe that for de Sitter with conformally invariant boundary conditions, this non-analyticity exists only when at least one field line (external or internal) is in the principal series (so that $\Delta\neq\Delta^{*}$).
@@ 6.3 Cosmological Optical Theorem for Flat FLRW
where we define each permutation symbol so that it is invariant under all real coordinate transforms in (time / space) except those for which $\det$ Jacobian =1=-1, in which case the (time / space) permutation symbol flips sign. [^48] If a Lorentz covariant parity-odd theory is desired, one simply requires the epsilon symbols to come together in the form of the usual spacetime permutation symbol: $\bm{\epsilon}^{a_{0}a_{1}\ldots a_{d}}=(d+1)\bm{\epsilon}^{[a_{0}}\bm{\epsilon}^{a_{1}a_{2}\ldots a_{d}]}$. However, for any analytic continuation that is continuously connected to the identity, the $\bm{\epsilon}$’s have to remain the same sign throughout, as the analytic continuation of 1 is 1. Now let a given amplitude be labelled $T=\pm 1$ according to whether it depends on $\bm{\epsilon}^{t}$ an (even/odd) number of times, and let it be labelled $X=\pm 1$ according to whether it depends on $\bm{\epsilon}^{i_{1}i_{2}\ldots i_{d}}$ an (even/odd) number of times.
@@
$\psi_{n}^{*}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};\Omega)=T\psi_{n}(\eta;e^{i\pi}\textbf{y}_{1},\dots,e^{i\pi}\textbf{y}_{n};e^{-i\pi}\Omega)\,$, (6.33)
@@
$\psi_{n}(\eta;-\textbf{y}_{1},\dots,-\textbf{y}_{n};\Omega)=X^{d}\psi_{n}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};\Omega)\,.$ (6.34)
@@
$\psi_{n}^{(L)}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};e^{i\pi}\Omega)=TX^{d}(-1)^{D(L-1)}\psi_{n}^{(L)}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};\Omega)=TX^{d}e^{i\pi D(L-1)}\psi_{n}^{(L)}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};\Omega)\,.$ (6.35)
@@
$\mathbf{RR}:\qquad\psi_{n}^{(L)*}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n})=e^{i\pi(d+1)(L-1)}\psi_{n}^{(L)}(\eta;-e^{i\pi}\textbf{y}_{1},\dots,-e^{i\pi}\textbf{y}_{n})\,.$ (6.36)
@@ 7.1 Unitarity and the Bunch-Davies Vacuum
$\textbf{k}\rightarrow e^{-i\theta}\textbf{k}\,.$ (7.11)
@@
This is a linear combination of sines and cosines and so could equivalently be represented as a shifted sine function which is non-zero at $\theta=0$. Therefore, this function is guaranteed to change sign over the range $0\leq\theta\leq\pi$ (or $\pi\leq\theta\leq 0$) and we must also rotate $\varepsilon$ alongside $\eta$, $\varepsilon\rightarrow\varepsilon e^{i\theta}$, so that the exponent is
@@ 7.2 The Loop Momentum Integral
where the prime in $\psi^{\prime(L)}_{n}$ is used to explicitly highlight that these wavefunction coefficients contain the $\delta$-function whilst those without the $\delta$-function will be denoted by $\psi_{n}^{(L)}$, and we have used the shorthand notation $\psi^{(L)}_{n}(\eta_{0};\textbf{k};\varepsilon)\equiv\psi_{n}^{(L)}(\eta;\textbf{k}_{1},\dots,\textbf{k}_{n};\epsilon)$. The external energies are indicated by $k_{a}=\lvert\textbf{k}_{a}\rvert$ whilst the $p_{i}$ are the internal energies which are built from the $\textbf{k}_{a}$ plus potentially some of the loop momenta, $\textbf{P}_{l}$. Note that here we have made the $\varepsilon$ prescription explicit and that there are no derivatives acting on any of our propagators. This is purely for the sake of notational simplicity and adding any derivatives will not add any technical complications.
@@
The first component we will look at is the time integral which runs from -$\infty$ to $\eta_{0}$ when we have a transformation involving the time coordinate it is these limits that pick up a minus sign. To remove this we redefine the variables inside the integral with $\eta_{v}\rightarrow-\eta_{v}$ [^52] Note that whether this minus sign is $e^{i\pi}$ or $e^{-i\pi}$ will be important but, as it can rotate in a different way depending on which transformation we are considering, we will keep things vague at this point. When we write explicit expressions for the transformations we will be careful to make clear in which direction the rotation goes.. The effect of this on the measure factor is then straightforward as everything is just some power of $\eta$.
@@
For positive $\lambda$ we can see that the scaling on each side agrees. Naively, on the left hand side continuing $\lambda\rightarrow-\lambda$ would appear to have no impact. However, from the right hand side we can see how our analytic continuation in the energies alters this conclusion, instead recovering a factor of $(-1)^{d}$. [^53] Now that we have identified how to analytically continue loop energies, this removes the barrier which prevented the authors of [28] from deriving the cosmological equivalent of cutting rules from generalised Unitarity (see [146] for a pedagogical review of flat space generalised Unitarity). To see this, we first note that the region, $\Gamma$, depends only on the squares of the energies and so is unchanged. However, each of our square roots of energies does pick up a factor of 1-1. As we now have a potentially non-integer exponent in the dimension of the integral $d$ we will need to be careful with $e^{i\pi d}$ vs $e^{-i\pi d}$. For $\bf{RR}$, in accordance with the discussion in Section 7.1, all rotations in the energy must be taken in the lower half plane and so $\lambda=\lambda e^{-i\pi}$. Therefore we have,
@@
$\lim_{\eta\rightarrow\eta_{0}}K_{k}(\eta)$ $=1\,$, $\lim_{\eta\rightarrow-\infty}K_{k}(\eta,\eta^{\prime})$ $\propto e^{ik\eta}\,$, (7.24)
$\lim_{\eta,\eta^{\prime}\rightarrow\eta_{0}}G_{k}(\eta,\eta^{\prime})$ $=0\,$, $\lim_{\eta,\eta^{\prime}\rightarrow-\infty}G_{k}(\eta,\eta^{\prime})$ $\propto e^{ik\eta}\,.$ (7.25)
@@
$\mathcal{O}_{\textbf{k}}^{*}(\eta)K_{k}^{*}(\eta)$ $=0,$ $\lim_{\eta\rightarrow\eta_{0}}K_{k}^{*}(\eta)$ $=1\,$, $\lim_{\eta\rightarrow-\infty}K_{k}^{*}(\eta)$ $\propto e^{-ik\eta}\,$, (7.29)
$\mathcal{O}_{e^{-i\pi}\textbf{k}}(\eta)K_{e^{-i\pi}k}(\eta)$ $=0,$ $\lim_{\eta\rightarrow\eta_{0}}K_{e^{-i\pi}k}(\eta)$ $=1\,$, $\lim_{\eta\rightarrow-\infty}K_{e^{-i\pi}k}(\eta)$ $\propto e^{-ik\eta}\,.$ (7.30)
@@
$-\mathcal{O}_{\textbf{k}}^{*}(\eta)G_{k}^{*}(\eta,\eta^{\prime})$ $=-i\delta(\eta-\eta^{\prime}),$ $\lim_{\eta\rightarrow\eta_{0}}-G_{k}^{*}(\eta,\eta^{\prime})$ $=0\,$, $\lim_{\eta\rightarrow-\infty}-G_{k}^{*}(\eta,\eta^{\prime})$ $\propto e^{-ik\eta}\,$, (7.32)
$\mathcal{O}_{e^{-i\pi}\textbf{k}}(\eta)G_{e^{-i\pi}k}(\eta,\eta^{\prime})$ $=-i\delta(\eta-\eta^{\prime}),$ $\lim_{\eta\rightarrow\eta_{0}}G_{e^{-i\pi}k}(\eta,\eta^{\prime})$ $=0\,$, $\lim_{\eta\rightarrow-\infty}G_{e^{-i\pi}k}(\eta,\eta^{\prime})$ $\propto e^{-ik\eta}\,$, (7.33)
@@
$\mathcal{O}_{\textbf{k}}^{*}(\eta)K_{k}^{*}(\eta)$ $=0,$ $\lim_{\eta\rightarrow\eta_{0}}K_{k}^{*}(\eta)$ $=1\,$, $\lim_{\eta\rightarrow-\infty}K_{k}^{*}(\eta)$ $\propto e^{-ik\eta}\,$, (7.35)
$\mathcal{O}_{\textbf{k}}(e^{-i\pi}\eta)K_{k}(e^{-i\pi}\eta)$ $=0,$ $\lim_{\eta\rightarrow\eta_{0}}K_{k}(e^{-i\pi}\eta)$ $=1\,$, $\lim_{\eta\rightarrow-\infty}K_{k}(e^{-i\pi}\eta)$ $\propto e^{-ik\eta}\,.$ (7.36)
@@
$\lim_{\eta\rightarrow\eta_{0}}G_{k}^{*}(\eta,\eta^{\prime})$ $=0\,$, $\lim_{\eta\rightarrow\eta_{0}}e^{i\pi(d+1)}G_{k}(e^{-i\pi}\eta,e^{-i\pi}\eta^{\prime})$ $=0\,$, (7.39)
@@
$\text{$\bf{RR}$}:\left[\psi^{\prime(L)}_{n}(\eta_{0};\textbf{k};\varepsilon)\right]^{*}$ (7.46)
@@
where in the penultimate line we have employed the Feynman-Euler relationship, VI+L=1V-I+L=1 and the rotation on k applies to all internal and external energies. Next we investigate Scale Invariance which rotates all momenta magnitudes as well as the conformal time,
@@
$\qquad\qquad\qquad\qquad\quad\times\prod_{a}^{n}K_{e^{\mp i\pi}k_{a}}(e^{\pm i\pi}\eta)\prod_{i}^{I}G_{e^{\mp i\pi}p_{i}}(e^{\pm i\pi}\eta,e^{\pm i\pi}\eta^{\prime})\delta^{d}\left(e^{\mp i\pi}\sum_{a}^{n}\textbf{k}_{a}\right)\,$ (7.52)
@@
$\text{$\bf{CRT}$}:\left[\psi^{\prime(L)}_{n}(\eta_{0};\textbf{k};\varepsilon)\right]^{*}$ (7.55)